OpenStudy (anonymous):

how do you find the matrix that depicts a shear, with a shear factor of 1?

OpenStudy (anonymous):

It is the identity matrix, but one of the zeros is changed to a one.

OpenStudy (anonymous):

so could you pls check this pdf: http://www.xtremepapers.me/CIE/Cambridge%20IGCSE/0581%20-%20Mathematics%20Extended/0581_w09_qp_4.pdf and see question 2 b ii ??

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 0 \\ 1 & 1\end{matrix}\right] * \left(\begin{matrix}x \\ y\end{matrix}\right)\] the bottom left is where you need to put your shear factor. I don't beleive one is it though because: x' = x and y' = y + λx where lambda is your factor. Try a negative number, or maybe I'm wrong.

OpenStudy (anonymous):

yeah, yeah, you're right. it is -1. thanks, A LOT :):)

OpenStudy (anonymous):

No problem